OpenStudy (anonymous):

the fraction25/100 can be reduced to5/20 having the same value by dividing both the numbers 25 and 100 by

OpenStudy (lgbasallote):

what is \[\huge \frac{25}{100} \div \frac {5}{20} \implies \frac{25 \div 5}{100 \div 20}\]

OpenStudy (lgbasallote):

@lavenders still there?

OpenStudy (anonymous):

im still here yes

OpenStudy (lgbasallote):

so what is the answer to what i just posted?

OpenStudy (anonymous):

25

OpenStudy (lgbasallote):

what is 25 divided by 5?

OpenStudy (anonymous):

\[\frac{25}{100}=\frac{5}{20}\frac{5}{5}\]Divide numerator and denominator by five to get the reduced form. Not that it isn't the completely reduced form, which would be 1/4.